OpenStudy (anonymous):

Write the equation of a line in slope-intercept form that has a slope of -2 and passes through the point (2, -8).

OpenStudy (anonymous):

y = mx + b x = 2 and y = -8 slope(m) = -2 now sub and what do you get ?

OpenStudy (anonymous):

you need to solve for b

OpenStudy (anonymous):

y = -2x - 4? @kelliegirl33

OpenStudy (pinknabastak):

y=-2x+4

OpenStudy (pinknabastak):

wait no its y=-2x-4

OpenStudy (pinknabastak):

@jr221 you are correct

OpenStudy (anonymous):

y = mx + b -8 = -2(2) + b -8 = -4 + b -8 + 4 = b -4 = b equation is : y = -2x - 4....you are correct

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

sure thing :)